OpenStudy (anonymous):

how many different ways can you choose three vowels out of the box, if there are 5 vowels in the box and others dont matter

OpenStudy (anonymous):

C(n,p) gives the number of ways of selcting p objects from n objects

OpenStudy (anonymous):

so here its C(5,3) = 5*4*3/(3!) =10

OpenStudy (ikram002p):

5!/2! *3!

OpenStudy (ikram002p):

and to continue as @matricked said c(n,p) is given by \[c(n,p)=\left(\begin{matrix}5 \\ 3\end{matrix}\right)=\dfrac{n!}{p!(n-p)!}\]

OpenStudy (ikram002p):

such as n!=1×2×3×4....×n